P rove that all real of the form a + b 2 , a; b ∈ Z form s a ring.


1
Expert's answer
2022-02-08T11:20:58-0500

Given, "a + b\sqrt2 , a; b ∈ Z"

Proof:

A ring is a set R equipped with two binary operations + (addition) and ⋅ (multiplication) satisfying the following three sets of axioms, called the ring axioms


1.R is an abelian group under addition, meaning that:

"(a+b)+c=a+(b+c)"

we have:

"(a_1+b_1\sqrt2+a_2+b_2\sqrt2)+a_3+b_3\sqrt2=a_1+b_1\sqrt2+(a_2+b_2\sqrt2+a_3+b_3\sqrt2)"

"a+b=b+a"

we have:

"a_1+b_1\sqrt2+a_2+b_2\sqrt2=a_2+b_2\sqrt2+a_1+b_1\sqrt2"


There is an element 0 in R such that a + 0 = a

we have:

element 0 in R: "a=b=0"

For each a in R there exists −a in R such that a + (−a) = 0

we have:

"a+b\sqrt2+(-a-b\sqrt2)=0"


2.R is a monoid under multiplication, meaning that:

"(a \u22c5 b) \u22c5 c = a \u22c5 (b \u22c5 c)"we have:

"((a_1+b_1\sqrt2)(a_2+b_2\sqrt2))(a_3+b_3\sqrt2)=(a_1+b_1\sqrt2)((a_2+b_2\sqrt2)(a_3+b_3\sqrt2))"


There is an element 1 in R such that a ⋅ 1 = a and 1 ⋅ a = a

we have:

element 1 in R: "a=1,b=0"


3.Multiplication is distributive with respect to addition, meaning that:

a ⋅ (b + c) = (a ⋅ b) + (a ⋅ c)

we have:

"(a_1+b_1\sqrt2)(a_2+b_2\sqrt2+a_3+b_3\sqrt2)=(a_1+b_1\sqrt2)(a_2+b_2\sqrt2)+"

"+(a_1+b_1\sqrt2)(a_3+b_3\sqrt2)"


(b + c) ⋅ a = (b ⋅ a) + (c ⋅ a)

we have:

"(a_2+b_2\sqrt2+a_3+b_3\sqrt2)(a_1+b_1\sqrt2)=(a_1+b_1\sqrt2)(a_2+b_2\sqrt2)+"

"+(a_1+b_1\sqrt2)(a_3+b_3\sqrt2)"